# SpecTool: A Benchmark for Characterizing Errors in Tool-Use LLMs

### v1

@@ 6 Metrics
GPT-4-0125-preview 0.71 0.84 1.00 0.94 1.00 0.93 0.94 0.94
xLAM-8x22b 0.68 0.87 0.96 0.92 1.00 0.94 0.91 0.93
@@
xLAM-8x7b 0.63 0.78 0.95 0.92 1.00 0.95 0.93 0.91
Code-Llama-13b 0.54 0.61 0.89 0.92 0.91 0.78 0.79 0.81
GPT-4o-turbo-2024-05-13 0.53 0.83 0.82 0.95 1.00 0.88 0.87 0.89
GPT-3.5-turbo-1106 0.53 0.67 0.84 0.89 0.99 0.79 0.77 0.81
@@
Meta-Llama3-8b 0.27 0.58 0.4 0.92 0.71 0.49 0.42 0.64
Mistral-7b-Instruct-v0.1 0.24 0.49 0.54 0.91 0.93 0.52 0.67 0.7
@@
Mixtral-8x7b-Instruct-v0.1 0.1 0.11 0.42 0.41 0.43 0.43 0.42 0.2
7 Experiments
We perform an in-depth assessment of widely used large language models, encompassing both proprietary, API-based systems and models with publicly accessible weights. This evaluation covers a broad range of models to examine performance, versatility, and the impact of architectural variations across both closed and open-source platforms.
We incorporate a one-shot in-context example along with task-specific instructions in our prompt setup. Our benchmarking includes a selection of high-performing proprietary models as well as open-weight models. For each open-weight model, we evaluate the chat-optimized version when available, unless otherwise indicated.
For models without a dedicated chat version, we apply standardized prompt templates tailored for publicly accessible versions. We ensure that each prompt includes precise format instructions to enable the model to generate API calls in a structured and parsable format. To account for server inconsistencies, we execute multiple rounds per API call, thus minimizing potential disruptions and ensuring consistent evaluation outcomes.
7.1 Results
One goal of our SpecTool benchmark is to quantify various error patterns critical for understanding model fallacies. For that, we’ve conducted a comprehensive assessment of several performant LLMs, including GPT-4 (Achiam et al., 2023), GPT-3.5-Turbo from OpenAI, Meta-Llama-3-8b (Dubey et al., 2024), Code-Llama-13b (Rozière et al., 2024), Vicuna-13b-16k (Chiang et al., 2023), Mistral and Mixtral models from Mistral AI (Jiang et al., 2023) (Jiang et al., 2024), and xLAM (Zhang et al., 2024). Interestingly, while these models all perform well on ToolBench (Guo et al., 2024), their failure patterns are different. Details of the error patterns discovered using SpecTool are shown in table 3. The table 3 qualitatively explains the how each error pattern is exhibited in common model outputs.
One discovery is that models specifically tuned to focus on API calling mechanisms are generally less likely to produce false information, often known as “hallucinating” when dealing with function names and argument names. On the contrary, models like Vicuna (Chiang et al., 2023) (Zheng et al., 2024) that are designed for chat-based interactions do have a higher rate of inaccuracies, particularly within incorrect argument names and function names. It suggests that the tuning process, specifically for API calls, can lead to a substantial reduction in these hallucination problems.
Our results indicate that both proprietary and open-weight models fine-tuned for API-calling tasks show strong performance in reducing ”Insufficient API Calls.” This metric suggests these models are particularly effective at selecting the correct tools from the available API options, demonstrating a high proficiency in expected tool utilization.
This research constitutes a parallel and complementary pursuit in tandem with the development of the xLAM model (Zhang et al., 2024). The techniques applied within this study have been successfully incorporated into the xLAM data pipeline, evidence of which can be discerned from the resultant table 4. Most striking is the superior performance of GPT-4, achieving the highest success rate amongst all models evaluated. This could potentially suggest that larger models harnessing the power of the scaling law and quality pretraining in the base model remain critical to the success of agent applications. Another noteworthy observation is the exemplary performance of the Code-Llamma-13b model (Rozière et al., 2024), surpassing many other generically-purposed models. Our hypothesis is that there may be some inherent similarity between coding-associated tasks and function calls. And that could be beneficial for models designed for coding purposes, thereby enhancing their performance on function-calling tasks.
Figure 3: Environment API similarity versus Model Accuracy. Each individual color represents a model and the shapes represent a unique environment. Each environment on avg contains 8-9 API Options.
@@ 7.1.1 Impact of Feedback on Model Performance
Figure 4: Comparison of Model accuracy with and without Feedback
7.2 Ablation Study
We performed an ablation study to assess the importance of the constructive feedback mechanism introduced in the SpecTool benchmark. This analysis aimed to highlight the contribution of the feedback process to the model’s decision-making capabilities. As shown in Figure 4, the feedback mechanism plays a pivotal role in enabling the model to refine its actions by concentrating on essential aspects, such as correctly identifying API names, selecting appropriate API arguments, and adhering to the expected data types for those arguments.
In scenarios where the feedback mechanism was absent, the model frequently repeated errors, such as incorrect function names, leading to a significant reduction in its task completion effectiveness. Without guidance, the model struggled to recover from initial mistakes, which compounded its challenges in achieving desired outcomes. Conversely, with the feedback mechanism in place, the model exhibited immediate improvements in subsequent interactions. For example, it was able to adjust argument types or incorrect/mismatched argument names efficiently in response to errors. These results underscore the critical role of feedback in enhancing the adaptability and accuracy of models, particularly in iterative and error-prone processes.
7.2.1 Impact of Irrelevance on Model Performance
Figure 5: Comparison of Incorrect Function Name Metric on Queries with and without Irrelevance
In a complementary analysis, we conducted an in-depth examination of a subset of queries drawn from two distinct API environments to investigate the propensity of models to generate fabricated API calls when presented with unexpected or irrelevant requests. In this context, irrelevance was specifically defined as the inclusion of an additional requirement or constraint that was entirely unrelated to the functionality or parameters supported by the given API environment. By doing so, we sought to measure the extent to which these irrelevant constraints influenced the model’s ability to generate accurate, relevant outputs versus its tendency to hallucinate fictitious API calls that were not part of the original environment.
Our findings show that most models tend to produce more errors related to incorrect function names, posing a challenge to deploying them effectively in real-world applications. This highlights the need for better handling of unexpected inputs to improve stability and accuracy across various requests. Figure 5 illustrates the performance decline on the subset of queries both in the presence and absence of irrelevant inputs, highlighting the impact of query relevance on model performance. In the figure we particularly focus on the accuracy drop for the metric IFN, hence showcasing how we can segregate model fallacies using this benchmark.
@@ 7.2.2 Impact of API Complexity on Model Performance
The following study investigated how structural similarities between different APIs affect model decision-making and accuracy across various environments. To conduct this analysis, we isolated unique, descriptive tokens in each API function name by removing tokens shared across all names within an environment. This approach enabled us to compute a more accurate similarity measure using the Jaccard metric, which reflects how structurally similar API names are within each environment. This metric is advantageous as it assigns higher similarity scores to structurally similar APIs, which may differ only slightly, while scoring dissimilar, functionally distinct APIs lower.
@@ 8 Conclusion
This paper introduced SpecTool , a benchmark for evaluating LLMs on tool-use tasks with a focus on error patterns and feedback. Our findings highlight the importance of task-specific fine-tuning, as models optimized for API calls perform more reliably and reduce hallucinations compared to chat-oriented models. Larger models, like GPT-4, demonstrate superior performance, affirming the significance of scaling and quality pretraining. Additionally, we showed that the feedback mechanism in SpecTool plays a crucial role in improving model accuracy, guiding models to correct their function-calling behavior. Overall, this study provides valuable insights for refining LLMs in tool-use applications.
@@
In this work, we recognize the importance of capturing the significant error patterns of the model, as addressing these failures in planning and function-calling can lead to long-term improvements in the model effectiveness, hence enhancing the reasoning capabilities of LLMs. By analyzing these metrics, developers can identify specific weaknesses in language models and focus on improvements in those key areas. This targeted approach can help enhance the overall performance and reliability of the models.
Our benchmark SpecTool is a continuously developed pipeline for better use cases. Presently, we have tested on tasks belonging to only a single family of actions, for example a task would be estimating weather details with API actions only related to weather. Future work involves analyzing more complex tasks with real-world queries and APIs, along with large-scale testing across different LLMs.

### v2

@@ 8 Results
GPT-4-0125-preview 0.73 0.84 1.00 0.94 1.00 0.93 0.94 0.94
xLAM-8x7B 0.72 0.79 0.95 0.92 1.00 0.95 0.93 0.91
GPT-4o-turbo-2024-05-13 0.7 0.59 0.93 0.88 0.94 0.92 0.88 0.92
GPT-3.5-turbo-1106 0.67 0.7 0.94 0.86 0.95 0.94 0.93 0.91
@@
Code-Llama-13B 0.41 0.43 0.69 0.7 0.66 0.72 0.7 0.68
@@
DeepSeek-R1-Distill-Qwen-14B 0.32 0.47 0.66 0.63 0.43 0.73 0.72 0.73
Meta-Llama3-8b 0.27 0.58 0.4 0.42 0.71 0.49 0.42 0.64
Mistral-7B-Instruct-v0.1 0.24 0.49 0.54 0.21 0.23 0.52 0.47 0.6
@@
DeepSeek-R1-Distill-Qwen-7B 0.11 0.19 0.31 0.35 0.11 0.23 0.37 0.37
Mixtral-8x7B-Instruct-v0.1 0.1 0.11 0.12 0.11 0.1 0.15 0.12 0.2
Table 4: Results of ToolScan Benchmark across several performant LLMs on the proposed error metrics. IAC, IAV, RAC, IFE, IAT, IAN, IFN are error metrics occurring in each API Calls. We show the success rate as well as the percentage of API calls not having the error metrics (Higher is Better).
8.1 Ablation Study
8.1.1 Impact of Irrelevance on Model Performance
Figure 3: Left: We see that percent of queries with Incorrect Function Name (IFN) error is higher when we augment queries with irrelevant terms (Higher is Better). Right: We see that model success rate is higher when model is provided with feedback which helps it to correct itself (Higher is Better).
We did an in-depth examination on how irrelevant constraints affect model accuracy in generating API calls across two distinct environments. Irrelevance is defined as additional, unrelated requirements conflicting with the given API environment. Our analysis measures the impact on success rate and the tendency to hallucinate fictitious API calls.
Results indicate that most models produce more errors in Incorrect Function Name (IFN), posing deployment challenges in the real-world scenario. This underscores the need for improved handling of unexpected inputs. Figure 3 (Left) highlights that the percentage of queries without IFN decreases when irrelevant constraints are introduced in the query. This underscores the importance of our benchmark in uncovering the exact reason for the failure of the models.
@@ 8.1.2 Impact of Feedback On Model Performance
We also performed an ablation study to assess the importance of the constructive feedback mechanism introduced in the ToolScan benchmark. As shown in Figure 3 (Right), the feedback mechanism plays a pivotal role in enabling the model to refine its actions by concentrating on essential aspects, such as correctly identifying API names, selecting appropriate API arguments, and adhering to the expected data types for those arguments.
In scenarios where the feedback mechanism was absent, the model frequently repeats errors, such as incorrect function names, leading to a significant reduction in its task completion effectiveness. Without guidance, the model struggled to recover from initial mistakes, which compounded its challenges in achieving desired outcomes. Conversely, with the feedback mechanism in place, the model exhibited immediate improvements in subsequent interactions. For example, it was able to adjust argument types or incorrect/mismatched argument names efficiently in response to errors. These results emphasize the critical role of feedback in enhancing the adaptability and accuracy of models, particularly in iterative and error-prone processes.
8.1.3 Impact of Output Format Method on Model Performance
Figure 4: Comparison of Structured Unstructured Format versus Incorrect Format Error (Higher is Better).
We investigated the impact of structured versus non-structured output formats on the performance of LLMs. A related study by (Tam et al., 2024) examined open-ended benchmarks such as (Cobbe et al., 2021) and (Fansi Tchango et al., 2022), evaluating the effects of various output formats—including loose string, JSON, and YAML—on model performance. Their findings indicate that structured formats negatively impact an LLM’s ability to generate and reason effectively. However, their study does not explicitly identify the underlying cause of this phenomenon. We hypothesize that this performance degradation is attributed to the increased token count required in structured formats, as additional spaces, colons, and indentation contribute to token inflation. This, in turn, reduces the proportion of meaningful tokens dedicated to the core task.
Our results in Figure 4 align with this observation, demonstrating a similar decline in performance due to Incorrect Format Errors (IFE) in our function-calling task. Notably, our task presents a higher level of complexity compared to the benchmarks in (Tam et al., 2024), as it involves a significantly larger API space with a greater variety of function calls, argument names, and argument types.
Figure 5: We see that environments which similar APIs tend to confuse the models leading to lower performance (Higher is Better).
@@ 8.1.4 Impact of API Complexity on Model Performance
The following study investigated how structural similarities between different APIs affect model decision-making and accuracy across various environments. To conduct this analysis, we isolated unique, descriptive tokens in each API function name by removing tokens shared across all names within an environment. This approach enabled us to compute a more accurate similarity measure using the Jaccard metric, which reflects how structurally similar API names are within each environment. This metric is advantageous as it assigns higher similarity scores to structurally similar APIs, which may differ only slightly, while scoring dissimilar, functionally distinct APIs lower. For details on how we remove shared tokens and compute the metric, refer to Appendix B.4.
@@ 9 Conclusion
This paper introduced ToolScan, a benchmark for evaluating LLMs on tool-use tasks with a focus on error patterns and feedback. Our findings highlight the importance of task-specific fine-tuning, as models optimized for API calls perform more reliably and reduce hallucinations compared to chat-oriented models. Larger models, like GPT-4, demonstrate superior performance, affirming the significance of scaling and quality pre-training. Additionally, we showed that the feedback mechanism, relevance and unstructured output format in ToolScan plays a crucial role in improving the model accuracy, guiding models to correct their function-calling behavior.
@@
In this work, we recognize the importance of capturing the significant error patterns of the model, as addressing these failures in planning and function-calling can lead to long-term improvements in the model effectiveness, hence enhancing the reasoning capabilities of LLMs. By analyzing these metrics, developers can identify specific weaknesses in language models and focus on improvements in those key areas. This targeted approach can help enhance the overall performance and reliability of the models. For future work we are going to design environments involving multiple family of actions to develop a more robust testing environment.
Appendix A Appendix
A.1 Prompts for Query Augmentation
In the following section, we list all the prompts used in the process of query generation.
A.1.1 Constraint Based Augmentation
Table LABEL:fig:table-5 here describes the constraint based query generation prompt with a sample few-shot examples from the spaceflight environment. For every new environment we have its specific set of few-shot examples.
A.1.2 Sentence Transform Based Augmentation
Table LABEL:fig:table-6 here describes the sentence transform based query generation prompt.
A.1.3 Irrelevance Based Augmentation
Table LABEL:fig:table-7 here describes the irrelevance based query generation prompt.
A.1.4 Query Data Verification
To generate high-quality queries from the proposed augmentation methods, we employ a systematic refinement process guided by multiple important criteria. Specifically, we generate the tool calls required to solve query by an LLM proficient in function-calling to evaluate and refine the generated queries. This process includes the following steps:
  • Feasibility Evaluation: We ensure that all requisites of the generated queries are solvable using the available API list. This involves verifying the compatibility of the queries with the functionalities provided by the APIs.
  • Constraint Validation: The quality of constraints is assessed by leveraging the APIs’ execution capabilities. We check whether the arguments specified in the queries are feasible, executable, and capable of yielding significant outputs.
  • Manual Review: To ensure diversity and relevance, we manually assess the difficulty level of each selected query. This step helps to curate a set of queries that are representative of various use cases and complexity levels.
By combining automated evaluation with manual review, we aim to produce a diverse set of high-quality queries that are both executable and meaningful in the context of the given API list.
A.2 Environments Selection for ToolScan
We curate different functions with their respective parameter requirements for each environment using the following open-source APIs:
  • Patent API
  • Spaceflight API
  • Movie API
  • Weather API
  • Other APIs
Appendix B Method
B.1 Query Selection for ToolScan
Based on the feasibility of solving the query given the provided APIs, we randomly select 150 base queries. We make sure that we don’t introduce any duplicate or closely related queries generated by the augmentation.
B.1.1 Queries Selected for Feedback Ablation Study
For the feedback vs no feedback ablation study we use the same 150 queries to analyse the impact of the specific feedback provided to the model for executing the query successfully. Within, the average model score the major impact is caused on incorrect argument name, incorrect argument type and incorrect function name.
B.1.2 Queries Selected for Irrelevance Ablation Study
For the irrelevance ablation study, we specifically gather queries from the irrelevance augmentation. We do this for the base queries from 3 environments (Spaceflight, Movie and Patent). In total, this ablation study is done using 60 queries.
B.1.3 Queries Selected for Output Format Ablation Study
For the JSON vs no JSON ablation study we use the same 150 queries to analyse the impact of the structured format vs unstructured format method provided to the model for executing the query successfully. We have only shown the impact on incorrect format errors, but there is also a drop in other error metrics as well.
B.1.4 Queries Selected for API Complexity Ablation Study
For this study, we randomly collect 10 queries per environment to maintain unanimous averaging of scores across all the provided environments. The queries are selected from out original 150 base queries.
B.2 Prompt for Benchmark Inference
Table LABEL:fig:table-8 describes a sample prompt used during the inference in the Benchmark.
B.3 Sample Trajectory of the Benchmark
Table 9 describes a sample trajectory and highlights over the specific feedback provided to execute the query.
B.4 Details on API Similarity Impact on Model Performance
For the API Similarity Calculation, lets take an example of Patent Enviornment with the following functions.
[get_patent_with_title,
get_patent_with_title_and_date,
get_patent_with_title_lte,
get_patent_with_title_neq,
get_patent_with_title_eq].
We remove the tokens common across the list of functions and convert the functions in the following way:
[patent_title,
patent_title_date,
patent_title_lte,
patent_title_neq,
patent_title_eq]
Post refinement we calculate the similarity of each function name w.r.t to another:
$J(f_{i},f_{j})=\frac{|f_{i}\cap f_{j}|}{|f_{i}\cup f_{j}|}$ (1)
for all unique pairs $(f_{i},f_{j})$, where $i<j$.
The overall average Jaccard similarity across all function name pairs is then given by:
$J_{\text{avg}}=\frac{1}{\binom{n}{2}}\sum_{1\leq i<j\leq n}J(f_{i},f_{j})$ (2)
where $\binom{n}{2}=\frac{n(n-1)}{2}$ represents the total number of unique function name pairs.
Table 5: Prompts for Constraint Based Augmentation
You are an expert in augmenting agent’s queries for given parameters. You will be given some few shot examples and a list of tools, with their descriptions and their required parameters and based on your knowledge you have to generate queries with constraints based on the tools provided.
Each of the answer is expected to follow the format below:
Query1: (constraint added query)
Query2: (constraint added query)
Queryn: ….
Your task is providing the queries with 3 or 4 constraints based on the list of tool calls available, using the following criteria:
1. Add additional constraints to the query but make sure the query has single fixed answer instead of a worded description.
2. Focus on making significant use of arguments of each tool call in an innovative way. Pay close attention on each Argument’s Description and their input requirements while using them in the query.
3. Answer does not hallucinate and is relevant to the given tools list.
4. Innovate around using functions that are similar but have a very distinct difference.
5. Correctly follow the provided format.
FEWSHOT EXAMPLES:
Query1: Find articles from NASA or SpaceNews related to Artemis missions, not including those from European Spaceflight, limited to 2 results published between January 1st, 2021, and December 31st, 2021.
Query2: Retrieve the most recent article related to SpaceX from Spaceflight Now and the corresponding report with a summary mentioning F̈alcon Heavy,̈ both published after January 1st, 2020.
Table 6: Prompts for Sentence Based Augmentation
You are an expert at refining an agent’s queries by rephrasing sentences for clarity and precision without altering their intended meaning.
Each of the answer is expected to follow the format below:
Query1: (sentence transformed query)
Query2: (sentence transformed query)
Queryn: ….
You will receive some queries along with a list of tools, including their descriptions and required parameters. You will also be provided with an example of original vs augmented query. Your task is to transform each query’s sentence structure, ensuring that the revised sentence preserves the same context and intent as the original. Follow these criteria to guide your rephrasing:
1. Maintain the query’s original context and aim, ensuring that the rephrased sentence yields the same single, definitive answer as the original.
2. Keep the rephrased sentence relevant to the tools provided, without introducing any additional information or hallucination.
3. Pay close attention to each tool’s specific parameters and descriptions, making sure the rephrased query aligns with the tools’ intended use.
4. Follow the provided format precisely for each rephrased query.
ONE-SHOT EXAMPLE:
Original Query: Retrieve two articles focused on the Artemis missions, sourced only from NASA or SpaceNews, ensuring that no content originates from European Spaceflight, and limit the publication date to between January 1, 2021, and December 31, 2021.
Augmented Query: Find articles from NASA or SpaceNews related to Artemis missions, not including those from European Spaceflight, limited to 2 results published between January 1st, 2021, and December 31st, 2021.
Table 7: Prompts for Irrelevance Augmentation
You are an expert at creating modified queries with intentional irrelevance while maintaining the core context of the tools and arguments provided.
Each of the answer is expected to follow the format below:
Query1: (irrelevance added query)
Query2: (irrelevance added query)
Queryn: ….
You will receive some queries along with a list of tools, including their descriptions and required parameters. You will also be provided with an example of original vs augmented query. Your task is to introduce irrelevance by creating queries that either:
1. Include a constraint or condition not relevant to any of the listed arguments, or
2. Request execution of one function but provide parameters meant for a different function.
3. Make a completely new request which is unrelated all the list of functions provided.
When modifying each query, follow these guidelines: 1. Ensure that the added irrelevance does not alter the core meaning but introduces a new element or mismatch that makes the query more challenging to interpret.
2. On the original query section, the response should not hallucinate; it should remain constrained to the parameters available, even if they don’t perfectly match.
3. Innovate by intentionally introducing subtle conflicts in argument requirements or irrelevant requests.
4. Maintain the provided format for each modified query.
ONE-SHOT EXAMPLE:
Original Query: Find articles from NASA or SpaceNews related to Artemis missions, not including those from European Spaceflight, limited to 2 results published between January 1st, 2021, and December 31st, 2021.
Augmented Query: Retrieve two articles specifically with summary title containing Artemis missions, exclusively from NASA or SpaceNews, excluding any pieces from European Spaceflight. Ensure the publication dates fall within January 1, 2021, to December 31, 2021. Set an offset of 4 for the search results.
Table 8: Sample Benchmark Execution Prompt
Based on the previous context and API request history, generate an API request or a response as an AI assistant. We detail the available tools with their respective name, description, input(parameters) of each action as follows:
    "name": "get_joke_of_the_day_by_category",
    "description": "Get the joke of the day of a specific category",
    "parameters": {
      "category": {
        "type": "string",
        "description": "Category of the joke from the jokes categories API.",
        "required": true
      "limit": {
        "type": "number",
        "description": "Number of jokes to output.",
        "required": false
    "name": "finish",
    "description": "Return an answer and finish the task.",
    "parameters": {
      "answer": {
        "type": ["string", "number", "array"],
        "description": "Finish task using the answer parameter.",
        "required": true
ONE-SHOT EXAMPLE:
Goal: Please List all the categories of jokes available and
provide me with 2 jokes of the day.
Trajectory:
Action: get_categories_of_joke with Action Input: {}
Observation: [’food’, ’politics’, ’art’, ’sports’]
Action: get_joke_of_the_day with Action Input: {limit: 2}
Observation: [{’id’ : 1, ’result’: A man staggers into an
emergency room with two black eyes and a five iron wrapped
tightly around his throat. Naturally the doctor asks him what
happened. } ...]
Your output should follow the provided format which specifies the specific function to call along with its respective input. The example format is as follows.
Action: [your action] with Action Input: [your action input] If no function call is required additionally, please conclude the response using the finish function.
Table 9: Sample Trajectory with Feedback
Goal: Could you give us a list of sports and a list of markets,
      starting from page 3 to see the variety of options?
Trajectory:
  Interaction 1:
    Action: list_of_markets_for_pinnacle_odds
    Action Input: {"sport_id": 1, "page_num": 3}
    Feedback: ERROR | Invalid argument name page_num for function
    list_of_markets_for_pinnacle_odds, Valid arguments
    are [’sport_id’, ’league_ids’, ’event_type’, ’event_ids’]
  Interaction 2:
    Action: list_of_sports_for_pinnacle_odds
    Action Input: {\"sport_id\": 1, \"page_num\": "3"}
    Feedback: ERROR | Invalid argument type for
        page_num for function list_of_sports_for_pinnacle_odds
        Valid argument type is ’int’.
  Interaction 3:
    Action: list_of_sports_for_pinnacle_odds
    Action Input: {\"sport_id\": 1, \"page_num\": 3}
    Observation: [{’id’: 1, ’p_id’: 29, ’name’: ’Soccer’,
            ’last’: 1688136047, ’special_last’: 1688136033,
            ’last_call’: 1688136047}, {’id’: 2, ’p_id’: 33,
            ’name’: ’Tennis’, ’last’: 1688136023,
            ’special_last’: 1688118591,
            ’last_call’: 1688136041}].....